# Testing the Testers: Human-Driven Quality Assessment of Voice AI Testing Platforms

### v1

@@ 1 Introduction
Apple’s March 2025 decision to delay Siri’s advanced features until 2026—with internal testing showing AI features like video summarization working only 67-80% of the time—illustrates a fundamental challenge facing the voice AI industry [cnbc2025siri, ainvest2025, gurman2025siri]. This reliability crisis affects an ecosystem of over 8.4 billion voice assistants globally [statista2024voice], with mission-critical deployments across healthcare (\$3.18 billion market by 2030) [grandview2025healthcare], finance (\$84.99 billion in banking AI spending projected by 2030) [statista2024banking], and customer service where 95% of interactions are expected to be AI-powered by 2025 [servion2024ai]. Testing constitutes the foundation of software reliability, yet the industry lacks any systematic framework to evaluate the quality of voice AI testing approaches themselves.
Voice AI testing presents extraordinary challenges beyond challenges for traditional software testing. Unlike deterministic systems where test outcomes are predictable (e.g., verifying that a sorting function $f$ always arranges [3,1,2] as ([1,2,3])), voice agents exhibit non-deterministic behavior where identical inputs produce varied responses. Voice AI testing must account for prosody variations, real-time turn-taking under 500ms, latency requirements, and accent diversity [chen2024prosody, gnani2024barge, retell2025latency]. Beyond these technical challenges, evaluation requires assessing subjective qualities—empathy, naturalness, appropriateness—alongside semantic understanding across infinite conversational paths. These factors make voice AI testing one of the most complex emerging quality assurance challenges nowadays. While we focus in this study on voice AI given its unique technical challenges, many of our framework’s principles apply to conversational AI more broadly, including text-based chatbots, which also exhibit volatility in outcomes due to their non-deterministic nature.
@@ 1.3 Methodological Approach
To answer these research questions, we developed a human-centered evaluation framework that measures both dimensions of testing quality through rigorous empirical methods. For simulation quality (addressing RQ1), we use pairwise human comparisons where evaluators assess which of two test conversations better meets specified criteria—an approach that yields higher inter-rater reliability than absolute scoring [thurstone]. We measure three core dimensions: Scenario Adherence (how well simulations follow test requirements), Human Naturalness (prosody, flow, and conversational coherence), and Persona Adherence (consistency with assigned behavioral traits). For evaluation quality, we first establish human ground truth through systematic assessment; next, we measure how accurately testing platforms align with human consensus across six metrics: customer satisfaction (CSAT), appropriate call closure, repetition avoidance, conversation progression, response consistency, and expected outcome achievement. This dual approach—with 10 human evaluators per comparison for simulations and per recording for evaluations—ensures robust, reproducible quality measurement.
To validate the framework and quantify performance gaps (addressing RQ2), we conducted a comprehensive empirical study leveraging three commercial platforms—Coval[coval], Cekura[cekura] and Evalion[evalion] —as Simulations and Evaluations providers.
@@ 2.1 Voice AI Evaluation Methodologies and Benchmarks
The evaluation of voice AI systems has evolved from single-metric approaches to comprehensive multi-dimensional frameworks. AudioBench [wang2025audiobench] represents the current state-of-the-art, introducing 8 distinct evaluation tasks across 26 datasets, with particular emphasis on the trinity of speech understanding, audio scene understanding, and paralinguistic voice features. Their use of LLaMA-3-70B-Instruct as a model-as-judge evaluator achieving highest correlation with human judgments presages the broader trend toward LLM-based evaluation. This comprehensive approach addresses limitations identified by Aksënova et al. [aksenova2021might], who demonstrated that traditional metrics like Word Error Rate fail to capture the full spectrum of speech recognition use cases and linguistic variation impacts.
The dialogue evaluation landscape has similarly matured through frameworks like DialogBench [ou2024dialogbench], which systematically evaluates LLMs across 12 dialogue tasks, and DynaEval [zhang2021dynaeval], which unified turn-level and dialogue-level evaluation using Graph Convolutional Networks. These advances respond to fundamental challenges identified by Khalid and Lee [khalid2022explaining], who exposed systematic biases in neural dialogue evaluation metrics through adversarial behavioral analysis. The URGENT Speech Enhancement Challenge [zhang2025urgent] further demonstrated the necessity of multi-dimensional evaluation, revealing language dependency issues in generative models that single metrics cannot capture.
@@ 2.2 Commercial Voice AI Testing Platform Landscape
The 2024-2025 period marks a watershed moment in the commercialization of voice AI testing, with multiple Y Combinator-backed startups addressing different aspects of the testing challenge. Coval [techcrunch2025coval], founded by former Waymo tech lead Brooke Hopkins with \$3.3M seed funding, pioneered the application of autonomous vehicle testing methodologies to conversational AI. Their simulation-first approach generates thousands of test scenarios from minimal inputs, supporting both voice and text interactions with customizable environments.
Cekura [cekura2024funding], serving 75+ customers across regulated industries with \$2.4M in funding, emphasizes automated scenario generation with custom personas emulating varied accents, background noise, and conversational styles. Their focus on hallucination detection and interruption rate metrics addresses quality dimensions critical for healthcare and financial services applications. Hamming AI’s [hamming2025platform] approach of using AI-driven “voice characters” for high-scale stress testing, capable of thousands of concurrent test calls, demonstrates the scalability requirements of modern voice AI systems. The comparative analysis by Leaping AI [leaping2024comparison] provides practitioners with the first systematic evaluation of these emerging platforms, though notably lacks the empirical validation our research provides.
@@ 2.3 Human-in-the-Loop Evaluation Challenges and Solutions
Smith et al. [smith2022human] fundamentally challenged the field with “Human Evaluation of Conversations is an Open Problem,” comparing five crowdworker-based evaluation methods and revealing substantial variability in inter-annotator agreement and statistical sensitivity. Their findings that single-model per-turn evaluation differs significantly from pairwise per-dialog assessment in both human agreement rates and ability to detect performance differences underscores the complexity of establishing ground truth for conversational quality.
The Chatbot Arena platform [zheng2023chatbot] offers a practical solution through large-scale crowdsourcing, collecting over 240K votes from 90K users across 100+ languages. Their implementation of Bradley-Terry models for ranking, achieving 72-83% agreement between crowd-users and experts, demonstrates that aggregated human judgment can provide stable quality signals despite individual variability. The platform’s anonymous battle format and Elo rating system elegantly address evaluation bias while maintaining statistical robustness [zheng2023judging].
@@ 2.4 LLM-as-a-Judge Evaluation Paradigm
The comprehensive survey by Li et al. [li2024llms] established the theoretical foundation for LLM-based evaluation, defining core components of evaluation type, criteria, evaluation items, and optional references that produce assessment scores, explanations, and actionable feedback. Empirical validation through MT-Bench and Chatbot Arena [zheng2023judging] demonstrated that GPT-4 achieves greater than 80% agreement with human evaluators, matching human-human agreement levels of 81%.
However, systematic biases constrain the reliability of LLM judges. Position bias manifests as up to 75% preference for first-positioned responses, verbosity bias leads to preference for longer responses regardless of quality, and self-enhancement bias creates 10-25% preference for self-generated content [wang2024when]. Mitigation strategies have shown promise: position swapping improves GPT-4 consistency from 65% to 77.5%, few-shot prompting provides evaluation examples that increase consistency, and reference-guided evaluation reduces failure rates from 70% to 15%.
@@ 2.5 Simulation-Based Testing and Synthetic Conversation Generation
Academic advances in dialogue simulation have produced increasingly sophisticated approaches to test data generation. The ChatChecker framework [mayr2024chatchecker] introduced non-cooperative user simulation specifically designed to expose system weaknesses, incorporating persona generation, user simulation, breakdown detection, and dialogue rating in a modular architecture that generalizes across dialogue systems without requiring reference dialogues.
Commercial implementations have focused on practical challenges of realistic conversation generation. The emergence of platforms like Roark [hamming2025blog], which transforms real customer interactions into automated test suites while preserving sentiment and timing characteristics, represents a production-first testing philosophy. These approaches recognize that conversations must end naturally—not simply when a model fails—representing sophisticated understanding of real-world dialogue dynamics.
@@ 3.1 Overview
For this experimental study, as a subject agent we leveraged an enterprise-ready inbound customer support agent provided by Sei Right[sei]. Sei is a leading provider of Voice AI Agents in the finance space. This agent handles real customer inquiries about loans, closings, and financial services. By testing against a production-grade agent rather than a demo system, we ensure the simulations and evaluations involved in testing their agent reflect the complexities inherent to voice AI deployment in the real-world, rather than a simplified academic example.
@@ 3.2.1 Simulation Quality: Conceptual Approach
In practice, each metric is mapped to a specific survey question. Human participants then compare paired simulations from different providers and select the superior option for each metric (with ties allowed). The rationale of using pairwise comparisons rather than relying on absolute ratings is that the latter suffer from subjective interpretation and low inter-rater reliability. This methodology is grounded in established psychometric theory [thurstone].
@@ 3.5.3 Ensuring High-Quality Crowdsourced Human Evaluations
All human evaluators participating in this study are recruited through Prolific [prolific2025]. For screening of participants, the eligibility criteria requires them to be U.S.-based, native English speakers, aged 18–50, and with at least a 90% approval rate on the platform. These criteria ensured evaluators possessed both cultural context for US-based customer service expectations and demonstrated reliability in crowdsourced tasks.
@@ 3.6.1 Simulation Quality Analysis
  • The League scoring system treats each comparison as a match, awarding one point for wins, zero for losses. This straightforward approach provides transparent, interpretable results that directly reflect the frequency of preference [pairwiseleague].
  • The Elo rating system, adapted from chess rankings, accounts not only for win-loss records but also for the relative strength of opponents. Starting from a baseline rating, platforms gain or lose points based on both the outcome and the expected difficulty of each comparison. This difficulty is measured on the difference of Elo scores of each opponent [elorating].
Additionally, in order to aggregate metrics together, we propose using Principal Component Analysis (PCA) to capture the maximum variance across correlated quality measures [pcaanalysis].
@@ 3.6.3 Statistical Significance Testing
We employ Cochran’s $Q$ test [cochran1950comparison] as our omnibus test to determine whether the three platforms differ significantly in their evaluation accuracy across all paired binary observations. Additionally, we conduct pairwise comparisons using McNemar’s test [mcnemar1947note] with Bonferroni correction [bonferroni1936teoria, dunn1961multiple] to control for multiple comparisons, setting the adjusted significance threshold at $\alpha=0.05/3=0.017$.
@@ 3.6.4 Effect Size Quantification
To quantify the practical significance of observed differences, we calculate Cohen’s $h$ [cohen1988statistical] for comparing proportions and Cohen’s kappa [cohen1960coefficient] for measuring agreement beyond chance. We translate effect sizes into real-world impact by expressing differences as additional correct evaluations per 1000 simulations, providing direct insight into practical implications.
@@ 3.6.5 Robustness Validation
We validate our findings through bootstrap confidence intervals (10,000 iterations) [efron1979bootstrap] and permutation tests [good2000permutation] to ensure results are not artifacts of specific methodological choices. Performance consistency is assessed through coefficient of variation analysis, and we verify convergence across multiple scoring system variants to confirm the stability of our conclusions.
@@ 4.3.1 Evaluation Setup
To ensure fair comparison across providers, we designed a controlled experimental framework where all three platforms execute identical test cases. Each platform received the same scenario descriptions and persona specifications, implemented as human-written prompts directly in their respective systems. All simulations for all providers were run against the same subject agent, an enterprise-ready inbound customer support agent provided by Sei Right [sei].
@@ 4.3.2 Methodological Approach: Pairwise Comparisons
We choose to use direct comparisons between audio files rather than independent ratings in isolation. The methodology is inspired by established evaluation practices in machine learning research. Platforms such as LMArena [lmarena] have popularized the use of pairwise comparisons to evaluate large language models, showing that head-to-head judgments yield clearer and more consistent rankings than scalar ratings. Similar findings have been observed in psychometrics and human-computer interaction research, where comparative judgments often lead to higher inter-rater agreement [thurstone].
@@ 4.5.1 Scoring systems
To achieve this, we applied two of the most widely used and well-established scoring methods for aggregating pairwise comparisons [pairwiseelo, pairwiseleague]:
@@
  • Elo system: Adapted from chess ratings, this method accounts not only for wins and losses but also for the relative strength of the competing simulations. This allows strong simulations to be rewarded more for beating other strong ones, while losses against weaker ones have a greater penalty [elorating].
@@ 4.5.2 Tie handling
  • Including ties: In the league system, a tie yields 0.5 points for each simulation. In the Elo system, ties are replaced by two artificial matches—one win for each system—in random order since Elo is not commutative [hamilton2024elo].
@@ 4.5.3 Metric Aggregation
  • PCA aggregation: Using Principal Component Analysis (PCA), the first component is taken as the score. This ensures the score captures the greatest variance across the metrics [pcaanalysis] in each dimension.
@@ 4.5.4 Normalization
To enable direct comparison, combination, and aggregation of all metrics, we normalized them to a common scale. Specifically, we applied min–max normalization [minmaxscaling] across providers and simulations, rescaling each metric to the range 0–100 for consistency and simplicity.
@@ 4.6.1 Participants
Each simulation pair was evaluated by $N_{s}=10$ participants recruited through Prolific [prolific2025]. To assess whether a larger participant pool was necessary, we conducted an additional analysis: from each survey we drew a random subset of 5 participants and recomputed the ranked scores. The resulting rankings were highly consistent with those obtained using all 10 participants, indicating that increasing the number of participants would not meaningfully alter the outcomes (see Section  4.7.2). On this basis, we concluded that surveying 10 participants per simulation pair was sufficient, while further scaling would only increase study costs without providing additional benefit. For context, a full survey with the $S_{s}=45$ simulations, $C=3$ providers and $N_{s}=10$ participants, has an approximate cost of USD 2,500.

### v2

@@ 1 Introduction
Apple’s March 2025 decision to delay Siri’s advanced features until 2026—with internal testing showing AI features like video summarization working only 67-80% of the time—illustrates a fundamental challenge facing the voice AI industry [10, 1, 25]. This reliability crisis affects an ecosystem of over 8.4 billion voice assistants globally [50], with mission-critical deployments across healthcare (\$3.18 billion market by 2030) [23], finance (\$84.99 billion in banking AI spending projected by 2030) [49], and customer service where 95% of interactions are expected to be AI-powered by 2025 [45]. Testing constitutes the foundation of software reliability, yet the industry lacks any systematic framework to evaluate the quality of voice AI testing approaches themselves.
Voice AI testing presents extraordinary challenges beyond challenges for traditional software testing. Unlike deterministic systems where test outcomes are predictable (e.g., verifying that a sorting function $f$ always arranges [3,1,2] as ([1,2,3])), voice agents exhibit non-deterministic behavior where identical inputs produce varied responses. Voice AI testing must account for prosody variations, real-time turn-taking under 500ms, latency requirements, and accent diversity [9, 21, 42]. Beyond these technical challenges, evaluation requires assessing subjective qualities—empathy, naturalness, appropriateness—alongside semantic understanding across infinite conversational paths. These factors make voice AI testing one of the most complex emerging quality assurance challenges nowadays. While we focus in this study on voice AI given its unique technical challenges, many of our framework’s principles apply to conversational AI more broadly, including text-based chatbots, which also exhibit volatility in outcomes due to their non-deterministic nature.
@@ 1.3 Methodological Approach
To answer these research questions, we developed a human-centered evaluation framework that measures both dimensions of testing quality through rigorous empirical methods. For simulation quality (addressing RQ1), we use pairwise human comparisons where evaluators assess which of two test conversations better meets specified criteria—an approach that yields higher inter-rater reliability than absolute scoring [51]. We measure three core dimensions: Scenario Adherence (how well simulations follow test requirements), Human Naturalness (prosody, flow, and conversational coherence), and Persona Adherence (consistency with assigned behavioral traits). For evaluation quality, we first establish human ground truth through systematic assessment; next, we measure how accurately testing platforms align with human consensus across six metrics: customer satisfaction (CSAT), appropriate call closure, repetition avoidance, conversation progression, response consistency, and expected outcome achievement. This dual approach—with 10 human evaluators per comparison for simulations and per recording for evaluations—ensures robust, reproducible quality measurement.
To validate the framework and quantify performance gaps (addressing RQ2), we conducted a comprehensive empirical study leveraging three commercial platforms—Coval[14], Cekura[7] and Evalion[18] —as Simulations and Evaluations providers.
@@ 2.1 Voice AI Evaluation Methodologies and Benchmarks
The evaluation of voice AI systems has evolved from single-metric approaches to comprehensive multi-dimensional frameworks. AudioBench [52] represents the current state-of-the-art, introducing 8 distinct evaluation tasks across 26 datasets, with particular emphasis on the trinity of speech understanding, audio scene understanding, and paralinguistic voice features. Their use of LLaMA-3-70B-Instruct as a model-as-judge evaluator achieving highest correlation with human judgments presages the broader trend toward LLM-based evaluation. This comprehensive approach addresses limitations identified by Aksënova et al. [2], who demonstrated that traditional metrics like Word Error Rate fail to capture the full spectrum of speech recognition use cases and linguistic variation impacts.
The dialogue evaluation landscape has similarly matured through frameworks like DialogBench [39], which systematically evaluates LLMs across 12 dialogue tasks, and DynaEval [55], which unified turn-level and dialogue-level evaluation using Graph Convolutional Networks. These advances respond to fundamental challenges identified by Khalid and Lee [32], who exposed systematic biases in neural dialogue evaluation metrics through adversarial behavioral analysis. The URGENT Speech Enhancement Challenge [56] further demonstrated the necessity of multi-dimensional evaluation, revealing language dependency issues in generative models that single metrics cannot capture.
@@ 2.2 Commercial Voice AI Testing Platform Landscape
The 2024-2025 period marks a watershed moment in the commercialization of voice AI testing, with multiple Y Combinator-backed startups addressing different aspects of the testing challenge. Coval [30], founded by former Waymo tech lead Brooke Hopkins with \$3.3M seed funding, pioneered the application of autonomous vehicle testing methodologies to conversational AI. Their simulation-first approach generates thousands of test scenarios from minimal inputs, supporting both voice and text interactions with customizable environments.
Cekura [6], serving 75+ customers across regulated industries with \$2.4M in funding, emphasizes automated scenario generation with custom personas emulating varied accents, background noise, and conversational styles. Their focus on hallucination detection and interruption rate metrics addresses quality dimensions critical for healthcare and financial services applications. Hamming AI’s [28] approach of using AI-driven “voice characters” for high-scale stress testing, capable of thousands of concurrent test calls, demonstrates the scalability requirements of modern voice AI systems. The comparative analysis by Leaping AI [33] provides practitioners with the first systematic evaluation of these emerging platforms, though notably lacks the empirical validation our research provides.
@@ 2.3 Human-in-the-Loop Evaluation Challenges and Solutions
Smith et al. [47] fundamentally challenged the field with “Human Evaluation of Conversations is an Open Problem,” comparing five crowdworker-based evaluation methods and revealing substantial variability in inter-annotator agreement and statistical sensitivity. Their findings that single-model per-turn evaluation differs significantly from pairwise per-dialog assessment in both human agreement rates and ability to detect performance differences underscores the complexity of establishing ground truth for conversational quality.
The Chatbot Arena platform [58] offers a practical solution through large-scale crowdsourcing, collecting over 240K votes from 90K users across 100+ languages. Their implementation of Bradley-Terry models for ranking, achieving 72-83% agreement between crowd-users and experts, demonstrates that aggregated human judgment can provide stable quality signals despite individual variability. The platform’s anonymous battle format and Elo rating system elegantly address evaluation bias while maintaining statistical robustness [57].
@@ 2.4 LLM-as-a-Judge Evaluation Paradigm
The comprehensive survey by Li et al. [34] established the theoretical foundation for LLM-based evaluation, defining core components of evaluation type, criteria, evaluation items, and optional references that produce assessment scores, explanations, and actionable feedback. Empirical validation through MT-Bench and Chatbot Arena [57] demonstrated that GPT-4 achieves greater than 80% agreement with human evaluators, matching human-human agreement levels of 81%.
However, systematic biases constrain the reliability of LLM judges. Position bias manifests as up to 75% preference for first-positioned responses, verbosity bias leads to preference for longer responses regardless of quality, and self-enhancement bias creates 10-25% preference for self-generated content [53]. Mitigation strategies have shown promise: position swapping improves GPT-4 consistency from 65% to 77.5%, few-shot prompting provides evaluation examples that increase consistency, and reference-guided evaluation reduces failure rates from 70% to 15%.
@@ 2.5 Simulation-Based Testing and Synthetic Conversation Generation
Academic advances in dialogue simulation have produced increasingly sophisticated approaches to test data generation. The ChatChecker framework [37] introduced non-cooperative user simulation specifically designed to expose system weaknesses, incorporating persona generation, user simulation, breakdown detection, and dialogue rating in a modular architecture that generalizes across dialogue systems without requiring reference dialogues.
Commercial implementations have focused on practical challenges of realistic conversation generation. The emergence of platforms like Roark [27], which transforms real customer interactions into automated test suites while preserving sentiment and timing characteristics, represents a production-first testing philosophy. These approaches recognize that conversations must end naturally—not simply when a model fails—representing sophisticated understanding of real-world dialogue dynamics.
@@ 3.1 Overview
For this experimental study, as a subject agent we leveraged an enterprise-ready inbound customer support agent provided by Sei Right[44]. Sei is a leading provider of Voice AI Agents in the finance space. This agent handles real customer inquiries about loans, closings, and financial services. By testing against a production-grade agent rather than a demo system, we ensure the simulations and evaluations involved in testing their agent reflect the complexities inherent to voice AI deployment in the real-world, rather than a simplified academic example.
@@ 3.2.1 Simulation Quality: Conceptual Approach
In practice, each metric is mapped to a specific survey question. Human participants then compare paired simulations from different providers and select the superior option for each metric (with ties allowed). The rationale of using pairwise comparisons rather than relying on absolute ratings is that the latter suffer from subjective interpretation and low inter-rater reliability. This methodology is grounded in established psychometric theory [51].
@@ 3.5.3 Ensuring High-Quality Crowdsourced Human Evaluations
All human evaluators participating in this study are recruited through Prolific [41]. For screening of participants, the eligibility criteria requires them to be U.S.-based, native English speakers, aged 18–50, and with at least a 90% approval rate on the platform. These criteria ensured evaluators possessed both cultural context for US-based customer service expectations and demonstrated reliability in crowdsourced tasks.
@@ 3.6.1 Simulation Quality Analysis
  • The League scoring system treats each comparison as a match, awarding one point for wins, zero for losses. This straightforward approach provides transparent, interpretable results that directly reflect the frequency of preference [46].
  • The Elo rating system, adapted from chess rankings, accounts not only for win-loss records but also for the relative strength of opponents. Starting from a baseline rating, platforms gain or lose points based on both the outcome and the expected difficulty of each comparison. This difficulty is measured on the difference of Elo scores of each opponent [17].
Additionally, in order to aggregate metrics together, we propose using Principal Component Analysis (PCA) to capture the maximum variance across correlated quality measures [20].
@@ 3.6.3 Statistical Significance Testing
We employ Cochran’s $Q$ test [11] as our omnibus test to determine whether the three platforms differ significantly in their evaluation accuracy across all paired binary observations. Additionally, we conduct pairwise comparisons using McNemar’s test [38] with Bonferroni correction [4, 15] to control for multiple comparisons, setting the adjusted significance threshold at $\alpha=0.05/3=0.017$.
@@ 3.6.4 Effect Size Quantification
To quantify the practical significance of observed differences, we calculate Cohen’s $h$ [13] for comparing proportions and Cohen’s kappa [12] for measuring agreement beyond chance. We translate effect sizes into real-world impact by expressing differences as additional correct evaluations per 1000 simulations, providing direct insight into practical implications.
@@ 3.6.5 Robustness Validation
We validate our findings through bootstrap confidence intervals (10,000 iterations) [16] and permutation tests [22] to ensure results are not artifacts of specific methodological choices. Performance consistency is assessed through coefficient of variation analysis, and we verify convergence across multiple scoring system variants to confirm the stability of our conclusions.
@@ 4.3.1 Evaluation Setup
To ensure fair comparison across providers, we designed a controlled experimental framework where all three platforms execute identical test cases. Each platform received the same scenario descriptions and persona specifications, implemented as human-written prompts directly in their respective systems. All simulations for all providers were run against the same subject agent, an enterprise-ready inbound customer support agent provided by Sei Right [44].
@@ 4.3.2 Methodological Approach: Pairwise Comparisons
We choose to use direct comparisons between audio files rather than independent ratings in isolation. The methodology is inspired by established evaluation practices in machine learning research. Platforms such as LMArena [36] have popularized the use of pairwise comparisons to evaluate large language models, showing that head-to-head judgments yield clearer and more consistent rankings than scalar ratings. Similar findings have been observed in psychometrics and human-computer interaction research, where comparative judgments often lead to higher inter-rater agreement [51].
@@ 4.5.1 Scoring systems
To achieve this, we applied two of the most widely used and well-established scoring methods for aggregating pairwise comparisons [43, 46]:
@@
  • Elo system: Adapted from chess ratings, this method accounts not only for wins and losses but also for the relative strength of the competing simulations. This allows strong simulations to be rewarded more for beating other strong ones, while losses against weaker ones have a greater penalty [17].
@@ 4.5.2 Tie handling
  • Including ties: In the league system, a tie yields 0.5 points for each simulation. In the Elo system, ties are replaced by two artificial matches—one win for each system—in random order since Elo is not commutative [26].
@@ 4.5.3 Metric Aggregation
  • PCA aggregation: Using Principal Component Analysis (PCA), the first component is taken as the score. This ensures the score captures the greatest variance across the metrics [20] in each dimension.
@@ 4.5.4 Normalization
To enable direct comparison, combination, and aggregation of all metrics, we normalized them to a common scale. Specifically, we applied min–max normalization [24] across providers and simulations, rescaling each metric to the range 0–100 for consistency and simplicity.
@@ 4.6.1 Participants
Each simulation pair was evaluated by $N_{s}=10$ participants recruited through Prolific [41]. To assess whether a larger participant pool was necessary, we conducted an additional analysis: from each survey we drew a random subset of 5 participants and recomputed the ranked scores. The resulting rankings were highly consistent with those obtained using all 10 participants, indicating that increasing the number of participants would not meaningfully alter the outcomes (see Section  4.7.2). On this basis, we concluded that surveying 10 participants per simulation pair was sufficient, while further scaling would only increase study costs without providing additional benefit. For context, a full survey with the $S_{s}=45$ simulations, $C=3$ providers and $N_{s}=10$ participants, has an approximate cost of USD 2,500.